### v2

@@ A.6 Proof of Theorem 3
Next, we check sender-optimality. Let $p=\bar{p}_{k}$. By Lemma 2.i, we have $y_{k}-px_{k}=y_{N}-px_{N}$. Suppose for a contradiction that there exists an equilibrium that the sender strictly prefers to the top-$k$ equilibrium. This equilibrium induces a point $(x,y)$ in $\mathcal{X}$ such that $x>x_{k}$ and $y-px\geq y_{N}-px_{N}=y_{k}-px_{k}$. After replacing $y$ with $y^{\ast}(x)$, these inequalities still hold, contrary to the claim.
Finally, we prove the claim. If $(x,y)$ is on the upper envelope of $\mathcal{X}$ and $x>x_{k}$, then by Lemma 3, there exists $k^{\ast}$ in $\{1,\ldots,k\}$ and $\lambda$ in $(0,1]$ such that
@@
where the second equality follows from (8) and the last inequality follows from Lemma 2.ii.
@@ A.7 Proof of Theorem 4
First, we compute an upper bound on the sender’s equilibrium utility, as a function of the price $p$. For each price $p$, let
@@
Moreover, we can further restrict the maximization in (9) to the upper envelope of $\mathcal{X}$, which is characterized in Lemma 3.
@@
At $p=\bar{p}_{N-1}$, the function $V$ is upper semicontinuous but not lower semicontinuous. At all other prices $p$, the function $V$ is continuous. Therefore, $V$ has a global maximizer, which must be in $[\bar{p}_{0},\bar{p}_{N}]$. We show that $V$ has a global maximizer in $\{\bar{p}_{0},\ldots,\bar{p}_{N}\}$. It suffices to show that for each $k=0,\ldots,N-1$, the function $V$ is convex over the interval $(\bar{p}_{k},\bar{p}_{k+1})$. Since $V$ is linear over $(\bar{p}_{N-1},\bar{p}_{N})$, it suffices to consider $k=0,\ldots,N-2$.
Fix $k\in\{0,\ldots,N-2\}$ and $p\in(\bar{p}_{k},\bar{p}_{k+1})$. By Lemma 2.i,
@@ A.8 Example for Theorem 4
which holds because $\pi_{1}+\pi_{2}\leq 1$. Therefore, either the top-11 11 equilibrium or the top-22 22 equilibrium is optimal. The top-11 11 equilibrium is optimal if $\pi_{1}\geq 2\pi_{2}$ and the top-22 22 equilibrium is optimal if $\pi_{1}\leq 2\pi_{2}$. Note that
@@ A.9 Proof of Theorem 5
Recall that the favorability order is defined in Section A.1. We begin with two statistical lemmas about this order.
@@
For completeness, we prove both lemmas in Appendix B.
For each $k$, the inequality $\underaccent{\bar}{p}_{k,n}<\bar{p}_{k,n}$ follows from Lemma 4 since $T_{n}^{k}=0$ if and only if $(T_{n}^{k},B_{n}^{N-k})=(0,0)$. The ordering of the price thresholds follows from Lemma 5. It remains to prove the equilibrium condition.
Fix $n,k\in[N-1]$. Under the top-$k$ strategy profile, it is clear that the sender is playing a best response to the receiver’s strategy. We characterize whether the receiver is playing a best response to the sender’s strategy. Suppose that the sender uses the top-$k$ messaging strategy. Further, suppose counterfactually that the receiver could check $n\wedge k$ of the recommended attributes and also $n\wedge(N-k)$ of the unrecommended attributes. By exchangeability, a sufficient statistic for the receiver’s observation is $(T_{n}^{k},B_{n}^{N-k})$. It follows from Lemma 4 that the conditional expectation
@@ A.10 Proof of Lemma 2
    00$\displaystyle$ 0 $\displaystyle=\sum_{j=0}^{N}\pi_{j}(q_{j}^{k}-q_{j}^{N})(v_{j}-\bar{p}_{k}),$
    00$\displaystyle$ 0 $\displaystyle<\sum_{j=0}^{N}\pi_{j}(q_{j}^{k+1}-q_{j}^{N})(v_{j}-\bar{p}_{k}).$
@@ B.1 Proof of Lemma 4
Thus, for $1\displaystyle 0\leq t\leq n\wedge k-$ 1 and $t+1\leq j\leq k-n\wedge k+t$, we have
@@
This ratio is weakly increasing in $j$. For $1\displaystyle 0\leq b\leq n\wedge(N-k)-$ 1 and $k+b+1\leq j\leq N-n\wedge(N-k)+b$, we have
@@ B.2 Proof of Lemma 5
    Thus, for $1\leq k\leq N$ and $1\displaystyle 0\leq j\leq k-n\wedge k-$ 1, we have
@@
    This ratio is weakly increasing in $k$, strictly so if $k\leq n$.
@@
    =$=\begin{cases}\frac{\binom{j\wedge k}{n\wedge k}\binom{(N-j)\wedge(N-k)}{n\wedge(N-k)}}{\binom{k}{n\wedge k}\binom{N-k}{n\wedge(N-k)}}&\text{if}~{}n\wedge k\leq j\leq N-n\wedge(N-k),\\ 0&\text{otherwise}.\end{cases}$

### v3

@@ A.6 Proof of Theorem 3
Next, we check sender-optimality. Let $p=\bar{p}_{k}$. By Lemma˜2.i, we have $y_{k}-px_{k}=y_{N}-px_{N}$. Suppose for a contradiction that there exists an equilibrium that the sender strictly prefers to the top-$k$ equilibrium. This equilibrium induces a point $(x,y)$ in $\mathcal{X}$ such that $x>x_{k}$ and $y-px\geq y_{N}-px_{N}=y_{k}-px_{k}$. After replacing $y$ with $y^{\ast}(x)$, these inequalities still hold, contrary to the claim.
Finally, we prove the claim. If $(x,y)$ is on the upper envelope of $\mathcal{X}$ and $x>x_{k}$, then by Lemma˜3, there exists $k^{\ast}$ in $\{1,\ldots,k\}$ and $\lambda$ in $(0,1]$ such that
@@
where the second equality follows from (8) and the last inequality follows from Lemma˜2.ii.
@@ A.7 Proof of Theorem 4
First, we compute an upper bound on the sender’s equilibrium utility, as a function of the price $p$. Let $\vee$ denote the maximum operator. For each price $p$, let
@@
Moreover, we can further restrict the maximization in (9) to the upper envelope of $\mathcal{X}$, which is characterized in Lemma˜3.
@@
The function $V$ is upper semicontinuous. Therefore, $V$ has a global maximizer, which must be in $[\bar{p}_{0},\bar{p}_{N}]$. We show that $V$ has a global maximizer in $\{\bar{p}_{0},\ldots,\bar{p}_{N}\}$. It suffices to show that for each $k=0,\ldots,N-1$, the function $V$ is convex over the interval $(\bar{p}_{k},\bar{p}_{k+1})$. Since $V$ is linear over $(\bar{p}_{N-1},\bar{p}_{N})$, it suffices to consider $k=0,\ldots,N-2$.
Fix $k\in\{0,\ldots,N-2\}$ and $p\in(\bar{p}_{k},\bar{p}_{k+1})$. By Lemma˜2.i,
@@ A.8 Example for Theorem 4
which holds because $\pi_{1}+\pi_{2}\leq 1$. Therefore, either the top-$1$ equilibrium or the top-$2$ equilibrium is optimal. The top-$1$ equilibrium is optimal if $\pi_{1}\geq 2\pi_{2}$ and the top-$2$ equilibrium is optimal if $\pi_{1}\leq 2\pi_{2}$. Note that
@@ A.9 Proof of Theorem 5
Recall that the favorability order is defined in Section˜A.1. We begin with two statistical lemmas about this order.
@@
For completeness, we prove both lemmas in Appendix˜B.
For each $k$, the inequality $\underaccent{\bar}{p}_{k,n}<\bar{p}_{k,n}$ follows from Lemma˜4 since $T_{n}^{k}=0$ if and only if $(T_{n}^{k},B_{n}^{N-k})=(0,0)$. The ordering of the price thresholds follows from Lemma˜5. It remains to prove the equilibrium condition.
Fix $n,k\in[N-1]$. Under the top-$k$ strategy profile, it is clear that the sender is playing a best response to the receiver’s strategy. We characterize whether the receiver is playing a best response to the sender’s strategy. Suppose that the sender uses the top-$k$ messaging strategy. Further, suppose counterfactually that the receiver could check $n\wedge k$ of the recommended attributes and also $n\wedge(N-k)$ of the unrecommended attributes. By exchangeability, a sufficient statistic for the receiver’s observation is $(T_{n}^{k},B_{n}^{N-k})$. It follows from Lemma˜4 that the conditional expectation
@@ A.10 Proof of Lemma 2
    $\displaystyle 0$ $\displaystyle=\sum_{j=0}^{N}\pi_{j}(q_{j}^{k}-q_{j}^{N})(v_{j}-\bar{p}_{k}),$
    $\displaystyle 0$ $\displaystyle<\sum_{j=0}^{N}\pi_{j}(q_{j}^{k+1}-q_{j}^{N})(v_{j}-\bar{p}_{k}).$
@@ B.1 Proof of Lemma 4
Thus, for $0\leq t\leq n\wedge k-1$ and $t+1\leq j\leq k-n\wedge k+t$, we have
@@
This ratio is weakly increasing in $j$. For $0\leq b\leq n\wedge(N-k)-1$ and $k+b+1\leq j\leq N-n\wedge(N-k)+b$, we have
@@ B.2 Proof of Lemma 5
    Thus, for $1\leq k\leq N$ and $0\leq j\leq k-n\wedge k-1$, we have
@@
    This ratio is strictly increasing $k$.
@@
[…]